# AutoTimes: Autoregressive Time Series Forecasters via Large Language Models

### v2

@@ Abstract
Foundation models of time series have not been fully developed due to the limited availability of time series corpora and the underexploration of scalable pre-training. Based on the similar sequential formulation of time series and natural language, increasing research demonstrates the feasibility of leveraging large language models (LLM) for time series. Nevertheless, the inherent autoregressive property and decoder-only architecture of LLMs have not been fully considered, resulting in insufficient utilization of LLM abilities. To further exploit the general-purpose token transition and multi-step generation ability of large language models, we propose AutoTimes to repurpose LLMs as Autoregressive Time series forecasters, which independently projects time series segments into the embedding space and autoregressively generates future predictions with arbitrary lengths. Compatible with any decoder-only LLMs, the consequent forecaster exhibits the flexibility of the lookback length and scalability of the LLM size. Further, we formulate time series as prompts, extending the context for prediction beyond the lookback window, termed in-context forecasting. By adopting textual timestamps as position embeddings, AutoTimes integrates multimodality for multivariate scenarios. Empirically, AutoTimes achieves state-of-the-art with $0.1\%$ trainable parameters and over 5×5$\times$5 × training/inference speedup compared to advanced LLM-based forecasters.
@@ 1 Introduction
Time series forecasting is of crucial demand in real-world applications, covering diverse domains including climate, economics, energy, operations, etc [22, 42]. The growing challenges of general-purpose forecasting, where one model is versatile to handle variable-length scenarios [24, 40] and the prediction is necessarily instructed by auxiliary information in other modalities [39, 43], underscores the demand for foundation models [3] of time series, which are aimed to exhibit enhanced capabilities, including multi-step generation, zero-shot generalization [48, 13], in-context learning and multimodal utilization [15], thereby expanding the scope of time series forecasting to a wider range of situations.
@@
Despite recent studies on large language models for time series (LLM4TS) achieving performance breakthroughs in current forecasting benchmarks [15], the mechanism by which LLMs are aligned to the time series modality still remains obscure. The pilot work, FPT [48] leverages LLMs as generic sequential representation extractors for time series, influencing subsequent LLM4TS methodologies. As depicted in Figure 1, the prevalent non-autoregressive approach, where time series are segmented into tokens, flattens and projects all lookback tokens for the prediction in a single step. However, it causes inconsistencies in both model structure and generative approach of LLMs: decoder-only models for autoregressive generation are adopted as encoder-only and non-autoregressive forecasters.
@@
Given that prior studies [9, 37] reveal that generalization capabilities of LLMs are largely derived from the decoder-only design trained autoregressively, talents of LLMs may not be fully exhibited. Further, autoregression, the essential concept of statistical forecasters like ARIMA [5] and exponential smoothing [38], empowers the flexibility on the prediction length of time series. Therefore, LLM-based forecasters can be potentially activated by the multi-step generation ability of LLMs, instantiated as one model for arbitrary forecast lengths with mitigated error accumulations [47].
@@
Figure 1: (a) Prevalent LLM4TS methods non-autoregressively generate predictions with the globally flattened representation of lookback tokens, while large language models essentially predict the next tokens by autoregression [46]. (b) Previous prompting mechanisms that concatenate language and time series tokens may lead to excessive context length and disparities in token modality.
@@
Motivated by the reflections, we propose AutoTimes to adapt LLMs as time series forecasters, which retrieves the consistency of autoregression with revitalized LLM capabilities to produce foundation models for time series forecasting. Technically, we independently embed time series segments into the latent space of the LLM by the consistent training objective, the next token prediction [2]. To fully leverage the inherent token transitions of LLMs and reduce the training cost, we freeze the LLM and train new embedding and projection layers of time series, which only account for up to $0.1\%$ total parameters. The consequent forecaster adopts autoregressive inference like LLMs, which is no longer constrained to specific lookback/forecast lengths. Going beyond conventional time series forecasting, we propose in-context forecasting as shown in Figure 3, where time series can be self-prompted by instructive downstream samples. We adopt LLM-embedded timestamps as the position embedding to utilize chronological information and align multiple variates. Our contributions lie in three aspects:
@@
  • By refining the inconsistency of non-autoregressive LLM4TS methods, we propose to inherit the autoregressive manner of LLMs, which frees our method from training respectively on the lookback length and allows arbitrary-length predictions with alleviated error accumulation.
@@
  • We present AutoTimes, a simple but effective approach to acquire LLM-based forecasters by lightweight adaptation, which utilizes the inherent token transition as the future extrapolation of time series. Further, we propose in-context forecasting, which renovates the conventional paradigm by introducing instructive time series prompts to enhance forecasting.
@@ 2.1 Autoregressive Models
Autoregression is an essential concept of both language modeling and time series forecasting. Despite prevalent deep forecasters [10, 26, 41, 47] adopt a non-autoregressive approach without the requirement of iterative forecasting, autoregression, the absent exploration in deep forecasters, serves as the fundamental principle of statistical methods, which enables variable-length predictions. The most well-known model, ARIMA [4] is developed by incorporating differencing on AR and MA models, which are both autoregressive models with learned time-invariant transition from the past to the future. Incorporated with decomposition and pre-defined transitions, exponential smoothing [38] and state space models (SSM) [12] also take the same autoregressive formulation.
Autoregressive language models [27, 31] are trained with fine-grained supervision, where the generated token of each position is independently supervised. Consequently, they are not constrained by specific input/output lengths and excel at multi-step generation. Furthermore, existing LLMs are inherently autoregressive models [46], which demonstrate advanced abilities that are not present in small models, such as the generalization [37], scalability [6], and task generality [31, 32]. Therefore, it is imperative to adapt off-the-shelf LLMs as autoregressive forecasters, which keeps the consistency to fully revitalize the model capacity and advanced capabilities.
@@ 2.2 Large Language Models for Time Series
With the immense advancement of large language model infrastructure, LLM4TS methods have been experiencing significant development in recent years. PromptCast [43] reformulates time series as text pairs and accomplishes forecasting as a sentence-to-sentence task. LLMTime [13] regards time series as numerical tokens, demonstrating the zero-shot generalizability in time series forecasting. FPT [48] fine-tunes parameters of LLMs to adapt it as a general representation extractor serving for multiple time series analysis tasks. UniTime [21] adapts a language model across diverse time series for a unified forecaster of multiple domains. Based on thriving prompting techniques, deft language prompts [15, 21] and soft prompting [7] for time series are further investigated.
LLM4TS methods have achieved great breakthroughs in the performance of time series forecasting, but the cost of training and inference can sometimes be unaffordable due to the immensity of LLMs. In contrast, AutoTimes introduces minimal parameters to obtain time series embeddings, keeping the LLM frozen to fully utilize the inherent token transition, thereby achieving better model efficiency. We further provide Table 1 to categorize existing LLM4TS methods by several essential aspects.
@@ 2.3 Multimodal Language Models
Multimodal models have been well-developed upon LLMs, among which vision language models (VLM) have experienced rapid growth [1, 27]. The booming pre-trained vision backbones [11, 30], together with the instruction tuning paradigm, has revealed the potential of LLMs for vision tasks, where visual tokens and language tokens are concatenated as the input of the LLM [19, 20]. Inspired by this, previous LLM4TS methods utilize instructive language tokens as prefix-prompts for time series analysis [15, 34, 43]. Unlike previous works, our proposed method regards time series itself as the instructive prompt. It avoids the modality gap caused by concatenating time series and language tokens directly. We further incorporate multimodal information, the textual timestamp of time series, such that the language model can effectively perceive readable date and periodicity [41] in a given context length, and align simultaneous events from different series [22] for multivariate forecasting.
Table 1: Comparison of LLM4TS methods: Autoregressive categories LLM-based forecasters by whether to conduct autoregression. Freeze LLM enables quick adaptation, which would otherwise require significant resources for fine-tuning. Multimodal refers to the utilization of other modalities. Prior to AutoTimes, none of the LLM4TS methods achieved all three.
@@ 3 Method
The proposed AutoTimes adapts large language models for multivariate time series forecasting. Given lookback observations $\mathbf{x}_{1:L}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{L}\}\in\mathbb{R}^{L\times C}$ with $L$ time steps and $C$ variates, the objective is to predict the future $F$ time steps $\mathbf{x}_{L+1:L+F}=\{\mathbf{x}_{L+1},\ldots,\mathbf{x}_{L+F}\}\in\mathbb{R}^{F\times C}$. Besides, the textual timestamp $\mathbf{a}_{t}$ (e.g. $2016/07/05$ $00$:$00$:$00$), as the most common covariate, is adopted for prediction, which is aligned with time points $\mathbf{x}_{t}\in\mathbb{R}^{C}$ at time $t$. The task is to train an LLM-based forecaster $f$ that is able to predict with the lookback length $L$ for the forecast length $F$ as:
@@
Figure 2: Overview of AutoTimes: (1) Time series and corresponding timestamps are segmented. (2) Textual timestamps are converted into the position embedding of segments. (3) AutoTimes learns to embed time series segments by next token prediction, where intermediate layers of LLM are frozen.
@@ 3.1 Modality Alignment
Time series segment
To empower the forecaster with the capability to predict the time series of arbitrary lengths, we reintroduce the autoregressive generation approach into time series forecasting. Prior to this, we define time series segment as the consecutive and non-overlapping time points of a single variate. It is regarded as the basic token of the LLM-based forecaster, which encompasses series variations and mitigates the excessive-step autoregression. To focus on modeling temporal variations, our forecaster predicts each variate independently. Beyond Channel Independence [26] that implicitly captures the multivariate correlation [22] by shared parameters, AutoTimes converts timestamps into position embeddings and explicitly aligns simultaneous segment tokens, which is detailed in the next paragraph. Therefore, we simplify $\mathbf{x}_{t}$ as the time point of specific variate $x_{t}\in\mathbb{R}$. Given a single-variate time series of context length $NS$, the $i$-th segment of length $S$ is denoted as:
@@
To fully reutilize the inherent token transition of large language models, we freeze the parameters of the LLM and drop the embedding and projection layers for the language token. Instead, we establish $\operatorname{SegmentEmbedding}(\cdot):\mathbb{R}^{S}\mapsto\mathbb{R}^{D}$, which independently embeds segments into the latent space:
@@
$\mathbf{TE}_{i}=\operatorname{SelectLast}\big{(}\operatorname{LLM}(\operatorname{TextualDuration}(\mathbf{s}_{i}))\big{)}.$ (4)
Notably, $\mathbf{TE}_{i}$ is pre-computed by LLMs such that runtime forwarding for language tokens is not required during training. Given that the latent space of the LLM locates both time series segments and language tokens, the position embedding can be integrated with the corresponding segment without increasing the context length. Concretely, the token embedding $\mathbf{E}_{i}\in\mathbb{R}^{D}$ is obtained by:
@@ 3.2 Next Token Prediction
Since prevalent LLMs [6, 35] are endowed with the capability of predicting the next token $\mathbf{s}_{i}$ based on the preceding tokens $\mathbf{s}_{<i}$ autoregressively, we reutilize LLMs in a fully consistent approach and accomplish the prediction of arbitrary lengths iteratively. Given a time series of context length $NS$, the input series is segmented and embedded into $N$ token embeddings $\{\mathbf{E}_{1},\dots,\mathbf{E}_{N}\}$. The training objective is to independently generate the next tokens $\{\hat{s}_{2},\dots,\hat{s}_{N+1}\}$. We feed the token embeddings $\mathbf{E}_{i}$ into the intermediate layers of the LLM, which inherently parameterize token transitions:
@@ 3.3 In-context Forecasting
In terms of time series forecasting, we propose to constitute the pair by lookback-forecast windows, which are exactly represented as successive time points from earlier historical observations. Hence, we use time series in target datasets as prompts, extending the context for prediction beyond lookback series. We denote the extended context as $\mathcal{C}$, which contains $m$ time series prompts $\operatorname{tsp}^{(j)}$:
$\mathcal{C}=\{\operatorname{tsp}^{(j)}\in\mathbb{R}^{L+F}\},\ j=1,\dots,m.$ (10)
During training, we first obtain an LLM-based forecaster on a source dataset and select time series prompts from the downstream target dataset based on a unified strategy. During inference, we ensure all the prompts appear before the window to be predicted, such that there is no data leakage from future information. As shown in Figure 3, we concatenate time series prompts with lookback series and feed them as the context of the forecaster, renovating previous forecasting paradigms as:
@@ 4 Experiments
We conduct thorough evaluations of the performance of AutoTimes, including time series forecasting, zero-shot forecasting, and the proposed in-context forecasting. Additional analyses are included to evaluate the generality, scaling behavior, and adaptation cost of large language models. Detailed code implementation for reproduction is provided in the supplementary material.

### v4

@@ Abstract
Foundation models of time series have not been fully developed due to the limited availability of time series corpora and the underexploration of scalable pre-training. Based on the similar sequential formulation of time series and natural language, increasing research demonstrates the feasibility of leveraging large language models (LLM) for time series. Nevertheless, the inherent autoregressive property and decoder-only architecture of LLMs have not been fully considered, resulting in insufficient utilization of LLM abilities. To fully revitalize the general-purpose token transition and multi-step generation capability of large language models, we propose AutoTimes to repurpose LLMs as Autoregressive Time series forecasters, which projects time series into the embedding space of language tokens and autoregressively generates future predictions with arbitrary lengths. Compatible with any decoder-only LLMs, the consequent forecaster exhibits the flexibility of the lookback length and scalability with larger LLMs. Further, we formulate time series as prompts, extending the context for prediction beyond the lookback window, termed in-context forecasting. By introducing LLM-embedded textual timestamps, AutoTimes can utilize chronological information to align multivariate time series. Empirically, AutoTimes achieves state-of-the-art with $0.1\%$ trainable parameters and over 5×5$\times$5 × training/inference speedup compared to advanced LLM-based forecasters. Code is available at this repository: https://github.com/thuml/AutoTimes.
@@ 1 Introduction
Time series forecasting is of crucial demand in real-world applications, covering various domains including climate, economics, energy, operations, etc. [22, 43]. The growing challenges of general-purpose forecasting, where one model is versatile to handle variable-length scenarios [24, 41] and the prediction is necessarily instructed by auxiliary information in other modalities [40, 44], underscore the demand for foundation models [3] of time series, which are aimed to exhibit enhanced capabilities, including multi-step generation, zero-shot generalization [49, 13], in-context learning and multimodal utilization [15], thereby expanding the scope of time series forecasting to a wider range of situations.
@@
Despite recent studies on large language models for time series (LLM4TS) achieving performance breakthroughs in current forecasting benchmarks [15], the mechanism by which LLMs are aligned to the time series modality still remains obscure. The pilot work, FPT [49] leverages LLMs as generic sequential representation extractors for time series, influencing subsequent LLM4TS methodologies. As depicted in Figure 1 (a), the non-autoregressive approach, where time series are segmented into tokens, flattens and projects all lookback tokens for the prediction in a single step. However, it causes inconsistencies in both model structure and generative approach of LLMs: decoder-only models for autoregressive generation are converted to encoder-only and non-autoregressive forecasters.
@@
Given that prior studies [9, 38] reveal that generalization performance of LLMs is largely derived from the decoder-only structure trained autoregressively, talents of LLMs may not be fully exhibited. It is also supported by the recent rethinking of previous LLM4TS methods [35], which generally lack the maintenance of autoregression, the essential characteristic of both large language models and statistical forecasters [5, 39]. Therefore, autoregressive LLM4TS methods are underexplored, which can potentially unlock multi-step generation like LLMs, presenting one model for arbitrary lengths.
@@
Figure 1: (a) Prevalent LLM4TS methods non-autoregressively generate predictions with the globally flattened representation of lookback series, while large language models inherently predict the next tokens by autoregression [47]. (b) Previous methods adopt language prompts that may lead to the modality disparity, while we find time series can be self-prompted, termed in-context forecasting.
@@
Motivated by the reflections, we propose AutoTimes to adapt LLMs as time series forecasters, which retrieves the consistency of autoregression with revitalized LLM capabilities to produce foundation models for time series forecasting. Technically, we independently embed time series segments into the latent space of language models by the consistent training objective: next token prediction [2]. To fully leverage the inherent token transitions of LLMs and reduce the training cost, we freeze the LLM and establish token embedding and projection for time series, which only account for up to $0.1\%$ total parameters. The consequent forecaster adopts autoregressive inference like LLMs, which is no longer constrained to specific lookback/forecast lengths. Going beyond conventional time series forecasting, we propose in-context forecasting as shown in Figure 1, where time series can be self-prompted by relevant contexts. We further adopt LLM-embedded timestamps as the position embedding to utilize chronological information and align multiple variates. Our contributions are summarized as follows:
@@
  • By refining the inconsistency of non-autoregressive LLM4TS methods, we propose to inherit the autoregressive property of LLMs, which frees our method from training respectively on the lookback length and allows arbitrary-length predictions with chronological awareness.
@@
  • We present AutoTimes, a simple but effective approach to acquire LLM-based forecasters by lightweight adaptation, which utilizes the inherent token transition as the future extrapolation of time series. Further, we propose in-context forecasting, which renovates the conventional paradigm by introducing relevant time series prompts to enhance forecasting.
@@ 2.1 Autoregressive Models
Autoregression is an essential concept of both language modeling and time series forecasting. Despite prevalent deep forecasters [10, 26, 42, 48] adopt a non-autoregressive approach without the requirement of iterative forecasting, autoregression, the absent exploration in deep forecasters, serves as the fundamental principle of statistical methods, which enables variable-length predictions. The most well-known model, ARIMA [4] is developed by incorporating differencing on AR and MA models, which are both autoregressive models with learned time-invariant transition from the past to the future. Incorporated with decomposition and pre-defined transitions, exponential smoothing [39] and state space models (SSM) [12, 23] also take the same autoregressive formulation.
Autoregressive language models [27, 31] are trained with fine-grained supervision, where the generated token of each position is independently supervised. Consequently, they are not constrained by specific input/output lengths and excel at multi-step generation. Furthermore, existing LLMs are inherently autoregressive models [47], which demonstrate advanced abilities that are not present in small models, such as the generalization [38], scalability [6], and task generality [31, 32]. Therefore, it is imperative to adapt off-the-shelf LLMs as autoregressive forecasters, which keeps the consistency to fully revitalize the model capacity and general-purpose token transitions.
@@ 2.2 Large Language Models for Time Series
With the immense advancement of large language model infrastructure, LLM4TS methods have been experiencing significant development in recent years. PromptCast [44] reformulates time series as text pairs and accomplishes forecasting as a sentence-to-sentence task. LLMTime [13] regards time series as numerical tokens, demonstrating the zero-shot generalizability in time series forecasting. FPT [49] fine-tunes parameters of the LLM to adapt it as a general representation extractor serving for multiple time series analysis tasks. UniTime [21] adapts a language model across diverse time series for a unified forecaster of multiple domains. Based on thriving prompting techniques, deft language prompts [15, 21] and soft prompting [7] for time series are further investigated.
LLM4TS methods have achieved performance breakthroughs in time series forecasting, but the cost of training and inference can sometimes be resource-consuming due to the immensity of LLMs. Recent revisiting of LLM4TS methods has revealed the inefficacy of LLMs adapted in the non-autoregressive approach [35]. By contrast, AutoTimes frozen LLMs, transfers the general-purpose token transition, and introduces minimal parameters to realize autoregressive next token prediction, thereby achieving better model efficiency and consistent utilization of large models. We further provide Table 1 that categorizes prevalent LLM4TS methods by several essential aspects.
@@ 2.3 Multimodal Language Models
Multimodal models have been well-developed upon LLMs, among which vision language models (VLM) have experienced rapid growth [1, 27]. The booming pre-trained vision backbones [11, 30], together with the instruction tuning paradigm, has revealed the potential of LLMs for vision tasks, where visual tokens and language tokens are concatenated as the input of the LLM [19, 20]. Inspired by this, previous LLM4TS methods utilize instructive language tokens as prefix-prompts for time series analysis [15, 34, 44]. Unlike previous works, our proposed method regards time series itself as the instructive prompt. It avoids the modality gap caused by concatenating time series and language tokens directly. We incorporate chronological information, the textual timestamp of time series, such that the language model can effectively perceive date and periodicity as the position embedding, and align simultaneous events from different time series [22] for multivariate forecasting.
Table 1: Comparison of LLM4TS methods: Autoregressive categories LLM-based forecasters by whether to conduct autoregression. Freeze LLM enables quick adaptation, which would otherwise require significant resources for fine-tuning. Multimodal refers to the utilization of information from other modalities. Prior to AutoTimes, none of the LLM4TS methods achieved all three.
@@ 3 Method
The proposed AutoTimes adapts large language models for multivariate time series forecasting. Given lookback observations $\mathbf{x}_{1:L}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{L}\}\in\mathbb{R}^{L\times C}$ with $L$ time steps and $C$ variates, the objective is to predict the future $F$ time steps $\mathbf{x}_{L+1:L+F}=\{\mathbf{x}_{L+1},\ldots,\mathbf{x}_{L+F}\}\in\mathbb{R}^{F\times C}$. Besides, the textual timestamp $\mathbf{a}_{t}$ (e.g. $2016/07/05$ $00$:$00$:$00$), as the most common covariate, is adopted for prediction, which is aligned with time points $\mathbf{x}_{t}\in\mathbb{R}^{C}$ at time $t$. The task is to train an LLM-based forecaster $f$ that is able to predict with the (varying) lookback length $L$ for the (arbitrary) forecast length $F$ as:
@@
Figure 2: An example to illustrate how AutoTimes adapts language models for time series forecasting.
@@ 3.1 Modality Alignment
Time series token
To empower the forecaster with the capability to predict time series for arbitrary lengths, we repurpose autoregressive LLMs as time series forecasters as depicted in Figure 2. Prior to this, we define time series token as the consecutive and non-overlapping segment of a single variate. It is regarded as the common token of the LLM-based forecaster, which encompasses series variations and mitigates excessively long autoregression. To focus on modeling temporal variations, our forecaster predicts each variate independently. Beyond Channel Independence [26] that implicitly captures the multivariate correlation [22] by shared parameters, AutoTimes converts timestamps into position embeddings and explicitly aligns simultaneous segment tokens, which is detailed in the next paragraph. Therefore, we simplify $\mathbf{x}_{t}$ as the time point of specific variate $x_{t}\in\mathbb{R}$. Given a single-variate time series of context length $NS$, the $i$-th segment of length $S$ is denoted as:
@@
Considering the general-purpose token transition, we freeze the parameters of large language models. To realize the token-wise alignment between time series tokens and language tokens, we establish $\operatorname{SegmentEmbedding}(\cdot):\mathbb{R}^{S}\mapsto\mathbb{R}^{D}$ that independently embeds segments into the latent space:
@@
$\mathbf{TE}_{i}=\operatorname{SelectLast}\big{(}\operatorname{LLM}(\operatorname{TimestampTemplate}(\mathbf{s}_{i}))\big{)}.$ (4)
Notably, $\mathbf{TE}_{i}$ is pre-computed by LLMs such that runtime forwarding for language tokens is not required during training. Given that the latent space of the LLM locates both time series tokens and language tokens, the position embedding can be integrated with the corresponding time span without increasing the context length. Concretely, the token embedding $\mathbf{E}_{i}\in\mathbb{R}^{D}$ is obtained by:
@@
Figure 3: Overview of AutoTimes: (1) time series and corresponding timestamps are segmented; (2) textual timestamps are converted into the position embeddings by the LLM; (3) time series segments are embedded and projected by next token prediction, where intermediate layers of LLM are frozen.
@@ 3.2 Next Token Prediction
As shown in Figure 3, prevalent LLMs [6, 36] are endowed with the capability of predicting the next token $\mathbf{s}_{i}$ based on the preceding tokens $\mathbf{s}_{<i}$. We reutilize LLMs in a fully consistent approach and generate prediction of arbitrary lengths iteratively. Given a time series of context length $NS$, the input series is segmented and embedded into $N$ token embeddings $\{\mathbf{E}_{1},\dots,\mathbf{E}_{N}\}$. The training objective is to independently generate the next tokens $\{\hat{s}_{2},\dots,\hat{s}_{N+1}\}$. We feed the token embeddings $\mathbf{E}_{i}$ into the intermediate layers of the LLM, which inherently parameterize token transitions:
@@ 3.3 In-Context Forecasting
In terms of time series forecasting, we propose to constitute the pair by lookback-forecast windows, which are exactly represented as successive time points from earlier historical observations. Hence, we use time series in target datasets as prompts, extending the context for prediction beyond consecutive lookback series. We denote the extended context as $\mathcal{C}$, which contains $m$ time series prompts $\operatorname{tsp}^{(j)}$:
$\mathcal{C}=\{\operatorname{tsp}^{(j)}=\mathbf{x}_{\leq t_{j}}|\text{ earlier historical time series}\},\ j=1,\dots,m,\ t_{j}\leq L.$ (10)
During training, we first obtain an LLM-based forecaster on a source dataset and select time series prompts from the downstream target dataset based on a unified strategy. During inference, we ensure all the prompts appear before the window to be predicted, such that there is no data leakage from future information. As shown in Figure 4, we concatenate time series prompts with lookback series and feed them as the context of the forecaster, termed in-context forecasting:
@@ 4 Experiments
We conduct thorough evaluations of the performance of AutoTimes, including time series forecasting, zero-shot forecasting, and the proposed in-context forecasting. Additional analyses are included to evaluate the generality, scaling behavior, and adaptation cost of large language models. Detailed code implementation for reproduction is provided in our public code repository.